# Measuring Star Formation Histories from Asymptotic Giant Branch Stars: A Demonstration in M31

### v2

@@ 5.3 Measuring the SFH of M31’s Inner Halo
We now provide an illustrative use case for our AGB star-only SFH technique by measuring the SFH for the inner stellar halo of M31. M31 is long-known to have a complex stellar morphology and a halo filled with substructure (see McConnachie et al. 2018 and references therein). Measuring the SFH of M31’s stellar halo from AGB stars may help shed new light on its formation history, while also providing a practical showcase for the power of AGB stars for measuring SFHs of stellar halos in other galaxies throughout the local Universe.
@@ 5.3.1 Data and Extinction Corrections
Our photometry includes data out to a projected radius of 32 kpc, encompassing M31’s inner stellar halo. We exclude data within a deprojected distance of 20 kpc, as shown in Figure 9, to avoid the complexity extinction in the disk (e.g., Dalcanton et al., 2015). Within the stellar halo footprint, we now verify the internal reddening derived from the extinction map discussed in §5.3 is very small and does not vary much spatially, using the the Draine et al. (2014) map which covers the full range of our UKIRT data in major axis, but not the full range in minor axis. However, the minor axis (northwest and southeast sections of M31’s halo) are expected to contain even less dust. The average extinction values derived from the corrected map discussed in §5.2.1 outside of a deprojected distance of 20 kpc are mag and mag, which are exactly equal to the foreground Milky Way reddening corrections also discussed in §5.2.1. For simplicity, we apply these small extinction corrections directly to our photometry before modeling the CMD, and then do not solve for extinction as part of the CMD modeling.
Figure 10 shows the NIR CMD of M31’s inner halo we used to measure its SFH ( kpc). The grey box indicates the selection region for AGB stars. Select PARSEC-COLIBRI isochrones (0.1, 1, 8 Gyr) are included for reference. There are AGB stars in this region. The 50% completeness limit across the inner stellar halo is mag.
@@ 5.3.2 Spatial Selection
The large number of AGB stars in M31’s inner halo allow us to measure a spatially resolved SFH. We divide our photometry into 18 spatial regions according to the following criteria. First, as illustrated in Figure 9, we prioritize placing the four major substructures from McConnachie et al. (2018) into separate regions. Regions 1 and 2 contain the Eastern and Western Shelves, respectively, which have been identified as being radial “shell” features created from the Giant Stellar Stream (GSS) progenitor’s last pericentric passage (Ferguson et al., 2002, 2005; Fardal et al., 2006, 2008; Richardson et al., 2008; Fardal et al., 2012; Bernard et al., 2015b). Region 3 contains NGC 205 (M110), a satellite galaxy of M31. Region 4 contains M32, another satellite galaxy of M31. D’Souza & Bell (2018) proposed that M32 is the progenitor of the GSS, whose merger with M31 is thought to be responsible for the burst of global star formation seen in M31’s disk and inner stellar halo Gyr ago. The remaining regions were chosen to have at least 700 AGB stars, which is near the lower limit for the number of AGB stars needed for a robust SFH recovery (see Appendix A). The regions were delineated via a combination of radial and azimuthal lines, resulting in the irregular region shapes seen in Figure 9. The goal for these remaining sections is simply to provide as much resolved spatially information as possible, particularly in the radial direction, in light of other spatial and number of star constraints. The number of AGB stars ranged from AGB stars in region 16 to AGB stars in region 3.
@@ 5.3.3 Star Formation History
We run MATCH on each of the 18 spatial sub-regions. In each instance, we adopt an age range of to dex with a spacing of 0.1 dex and a metallicity range of [M/H] to 0.0 with a spacing of 0.2 dex. We compute random uncertainties for each region following using the HMC approach described in §3.
Figure 11 and Figure 12 show the differential and cumulative SFHs for each region. We note that there is fairly sizable variation in timing, duration, and amplitude of the SFHs as a function of position. Because the inner stellar halo of M31 is dominated by sub-structure, and not a clear smooth component, it is perhaps not surprising to see such variations in the SFHs and stellar mass growth as the inner halo was assembled from many individual pieces, each with their own formation histories.
In general, we find that most regions have prominent star formation from Gyr ago and again Gyr ago. Many also exhibit some level of star formation at ages of Gyr. In terms of cumulative stellar mass growth, there is a factor of difference in the mass formed in the least (region 13) and most (region 3) active regions. Region 3 contains NGC 205, while region 13 contains among the least amount of substructure according to the maps in McConnachie et al. (2018). It may be the closest we come to sampling the smooth underlying halo of M31. Other sparse-looking regions (e.g., 14 and 16) are located in the direction of the GSS. Expanding the NIR imaging and/or applying this method to other wide-field NIR imaging of M31 is an obvious path forward to exploring the SFH of its stellar halo at larger radii and to providing broader context with which to interpret our SFHs.
In Figure 13, we plot the spatially-resolved SFH of M31’s inner stellar halo. Each of the 18 regions is color-coded by the best fit SFR at select lookback times. We note several interesting features. First, the most active regions are 3 and 8. Region 3 (NGC 205) shows prominent star formation and 2 Gyr ago, while region 8 also shows activity at and Gyr ago. Region 16, which is in the direction of the GSS, has its highest activity levels from Gyr ago. An inner “ring” of regions ( kpc) all show enhanced activity Gyr ago. Interestingly, we see little evidence of star formation in most regions at ages Gyr ago, aside from around NGC 205, and very little activity at ages Gyr ago.
Figure 14 shows the global SFH of M31’s inner halo (i.e., all 18 regions summed together), along with the corresponding 68% uncertainties. The overall shape of the SFH is similar to many of the individual regions, with modest star formation Gyr ago, a prominent burst Gyr ago, and steady and variable star formation from Gyr ago. Our global SFH shows that over the past Gyr, the inner stellar halo of M31 formed , % of which appears to have occurred Gyr ago.
@@
Figure 13 shows clearest enhancements in SFR Gyr ago. The highest SFRs at this time include the present day location of NGC 205 (region 3) and the the area adjacent to it that extends to the south (region 8). The SFH of NGC 205 measured from deep HST optical imaging shows a similar enhancement in star formation activity Gyr ago (Savino et al. in prep). Older studies of NGC 205’s stellar populations qualitatively support the presence of intermediate age and even young stellar populations in NGC 205’s central region (e.g., Richer et al., 1984; Lee, 1996; Davidge, 2003, 2005; Rose et al., 2005; Sharina et al., 2006; Monaco et al., 2009). To the best of our knowledge, there are no other SFHs overlapping region 8 in the literature.
@@
The interaction history of NGC 205 and M31 is not as well-studied as others in the M31 system, such as M32. There is a putative tidal ‘arc’ of NGC 205 that extends to the north (regions 3, 10, 13; McConnachie et al. 2004), though its association with NGC 205 is not definitive (e.g., Howley et al., 2008; Liu et al., 2024). The SFHs of NGC 205 and the arc regions appear to be anti-correlated. The similarity of SFHs between NGC 205 (region 3) and region 8, may suggest an association between the two (e.g., region 8 could be material stripped from NGC 205), but the lack of obvious tidal features in that direction from ground-based data (e.g., in the PAndAS stellar maps, Richardson et al. 2011; McConnachie et al. 2018) does not support this suggestion. An interesting future angle would be to combine wide-area kinematics, such as measured from recent work with DESI (Dey et al., 2023), and potentially Subaru/PFS in the future, with the AGB-star based SFHs to see if there are clear relationships between ages and kinematics, analogs to the disk heating study of Dorman et al. (2015). Finally, NGC 205 currently has no published proper motions. An orbital history, in tandem with our SFHs, would be useful in helping to decipher the interaction history between NGC 205 and M31.
Another prominent sub-structure in the vicinity of M31 is M32, which is located in region 4. Our SFH of region shows a prominent burst from Gyr ago and lower levels from Gyr ago. This is qualitatively similar to the SFH measured from moderately deep HST/ACS imaging (Savino et al. in prep). Monachesi et al. (2012)’s SFHs from two HST/ACS HRC fields that reach the oldest MSTO in M32 show broadly similar trends, however their coarse time resolution (i.e., 2-5 Gyr, 5-12 Gyr bins) make direct comparison challenging. Several regions adjacent to M32 (regions 6, 7, 11, 15) also show enhanced star formation Gyr, which is in line with some suggestions of when M32 (or its progenitor) may have started to interact with M31. The literature is rich with plausible interaction scenarios between M32 and M31 (e.g., Cepa & Beckman, 1988; Bekki et al., 2001; Choi et al., 2002; Ibata et al., 2004; Block et al., 2006; Gordon et al., 2006; Dierickx et al., 2014; D’Souza & Bell, 2018). Further comparisons with predictions from detailed simulations and our SFHs may help support or rule out some of these possibilities. A detailed exploration of these scenarios is beyond the intended scope of this paper. As with NGC 205, a proper motion and orbital history of M32 is likely to greatly help interpretation.
@@
Figure 15 shows the ‘quenching’ time in each region. As a proxy for ‘quenching’ we use the lookback time at which 90% of the stellar mass formed in each region’s cumulative SFH. Regions with larger values of stopped forming stars before regions with smaller values of .
In Figure 15, M32 and adjacent regions 4, 6, 7, and 11 show similar values of Gyr ago, which may suggest related stellar populations. Similarly, NGC 205 and adjacent regions 5, 8, and 17 all have similar values of Gyr ago. Regions 1 and 2, which contain the Eastern and Western shelves, respectively, have among the youngest stellar populations in the entire inner halo with Gyr ago. Finally, region 16, which is in the direction of the GSS is slightly younger, Gyr, than regions containing shelves.
On aggregate, this map indicates that star formation shut down closer to the disk of M31 Gyr prior to slightly more distant regions. However, the limitations of our spatial coverage, i.e., no extended and/or smooth stellar halo, make in challenging to interpret this result in a broader context. For example, the M31-M32 interaction scenario proposed by D’Souza & Bell (2018) indicates that the youngest stars should be more centrally concentrated than older stars. However, they considered a much larger spatial area overall, and a more targeted comparison is needed to see if our data are truly in tension with this scenario.
@@ 6 Caveats & Limitations
First, in this paper, we only examined AGB star-based SFHs back to Gyr ago. This oldest age restriction was motivated by the conservative choice to empirically vet our data against another SFH. However, in principle, as discussed in §2, it is possible to measure an AGB star-based SFH back to the oldest epoch of star formation (i.e., Gyr ago). While these ancient stars are unlikely to be found in the TP-AGB phase, they will appear as E-AGB stars on a NIR CMD. Indeed, ancient MW GC (e.g., NGC 6752; Norris et al. 1981) do have confirmed AGB stars, providing affirmation of this possibility. The limiting factor for testing this possibility is the lack of a suitable dataset (i.e., it must include the ancient MSTO and a suitable number of AGB stars). We plan to continue this investigation in future papers.
Second, as illustrated in Appendix A, a robust SFH requires several hundred AGB stars. Though our tests show that it is possible to recover a reliable SFH from as few as AGB stars, we adopt a more conservative recommendations of AGB, as a means of guarding against different types of SFHs, lower quality of data, etc. that are not conisdred in the present analysis.
@@
This minimum number is comfortably met by systems into the dwarf galaxy regime. For example, the SMC () and NGC 6822 (; McConnachie 2012) each contain AGB stars (e.g., Cioni & Habing, 2005; Boyer et al., 2011), while other fainter systems have several hundred AGB stars as well (e.g., Dalcanton et al., 2012a; Boyer et al., 2015). While galaxy luminosity provides a good rule of thumb for how many AGB stars to expect, SFH is clearly an important component. Systems that are purely ancient will have fewer AGB than those with more recent star formation (i.e., within the last several Gyr). Even fairly luminous systems with ‘pauses’ in the SFHs at intermediate ages (e.g., WLM; Albers et al., 2019; McQuinn et al., 2024a) may have fewer AGB stars than their luminosities would suggest.
@@
Third, our AGB star-only SFHs have no systematic uncertainties. Dolphin (2012) presents a means to estimate uncertainties on SFHs measured from resolved star CMDs that are designed to capture variations in the SFHs due to choice in stellar models. The method for computing these systematics relies on being able to measure SFHs on test datasets using different stellar models. Because COLIBRI uniquely provide suitable AGB star models, it is not possible to compute analogous systematic uncertainties for our AGB star-only SFHs. Even within the COLIBRI models, specific choices are made that can affect the age characteristics of the AGB stars. For example, our basis models for this analysis were generated with a Reimers mass loss value of for RGB stars. Large changes to this value can affect the evolution of AGB stars. For example, a much higher mass-loss rate on the RGB does not produce AGB stars at the oldest ages. It is not clear that such large mass-loss rates on the RGB are supported by data, but we nevertheless note this possibility. This type of choice is not unique to AGB star models. For example, the modeling of metal-diffusion in lower-mass stars can affect the lifetimes of MSTO stars, which can lead to systematically younger or older ages (e.g., Hidalgo et al., 2018).
@@
Finally, we note that only NIR data is suitable for AGB star-only SFHs. In the optical, there is too much overlap between the RGB and AGB to provide robust SFHs. In limited exploration, this appears to be true even for a single optical (e.g., F814W or F090W) and NIR band. We have not conducted an exhaustive exploration of all possible filter combinations, but our findings thus far strongly suggest this approach is optimal with two NIR filters. Age sensitivity is improved when both sides of the typical AGB SED peak are comfortably sampled (e.g., one band at m and another at m).
@@ 7 Future Prospects
The type of JWST dataset needed to measure AGB-star SFHs is well-matched to the JAGB (e.g., Lee et al. 2024; Li et al. 2024), and to some extent, TRGB distance programs (e.g., Anand et al. 2024a, b; Hoyt et al. 2024; Newman et al. 2024). The JAGB method requires at least a few hundred AGB stars in the J region of the NIR CMD and it often measured in filter combinations that sample either side for the SED peak of AGB stars. The high luminosity of AGB stars coupled with the characteristics of JWST suggest that AGB-based SFHs and JAGB distances already work out to 20 Mpc (e.g., Anand et al., 2024a, b; Hoyt et al., 2024; Lee et al., 2024) using hours of integration time. It is seems clear that JAGB distances and our SFH method can be extended to galaxies out to many tens of Mpc with suitably well-planned observations (e.g., to avoid too much crowding).
Programs that aim to measure TRGB distances with JWST may also be suitable for AGB-based SFHs. The filter combination provides for precise TRGBs, but is not particularly good for clearly separating all AGB and RGB stars in CMD space (see e.g., various CMDs in Anand et al. 2024a, b; Newman et al. 2024). However, the filter combination appears to enable precise (.5%) TRGB distances (Newman et al., 2024) and appears to be a good filter combination of AGB star-based SFHs as AGB stars are both brighter in these bands, have little overlap with the RGB, and sample both sides of the SEDs peak for most AGB stars. We illustrate this further in Appendix C.
Roman and Euclid offer similar potentials for AGB star-based SFHs, though their filter ranges are a little more limited than JWST. We have not explored the specific capabilities of these facilities, but based on the analysis in this paper, AGB star SFHs from their NIR CMDs appear promising as well.
@@
However, robust determinations of SFHs derived from AGB stars will require investigations into the full systematic uncertainties of the fits, similar to what has been done for SFHs based on optical CMDs (e.g., Weisz et al., 2011; Dolphin, 2012). As discussed in §6, we highly advocate for the development of an independent set of very detailed TP-AGB models. In addition to aiding AGB SFH studies like this one and others described in §1, another set of AGB models would improve our knowledge of the contribution of AGB stars to the integrated SEDs of galaxies at all cosmic epochs (e.g., Melbourne et al., 2012) and help with the calibration of the AGB star extragalactic distance scale, which use the P-L relations of Mira variable stars (e.g., Huang et al. 2020, 2024) or the mean luminosities of carbon stars (JAGB method; e.g., Madore & Freedman 2020; Ripoche et al. 2020; Zgirski et al. 2021; Lee et al. 2024).

### v3

@@ 5.3 Measuring the SFH of M31’s Inner Halo
We now provide an illustrative use case for our AGB star SFH technique by measuring the SFH for the inner stellar halo of M31 in 6 spatially-separated regions. M31 is long-known to have a complex stellar morphology and a halo filled with substructure (see McConnachie et al. 2018 and references therein). Measuring the SFH of M31’s stellar halo from AGB stars may help shed new light on its formation history, while also providing a practical showcase for the power of AGB stars for measuring SFHs of stellar halos in other galaxies throughout the local Universe.
@@ 5.3.1 Data and Extinction Corrections
Our photometry includes data out to a projected radius of 32 kpc, encompassing M31’s inner stellar halo. We exclude data within a deprojected distance of 20 kpc, as shown in Figure 9, to avoid the complexity extinction in the disk (e.g., Dalcanton et al., 2015). Within the stellar halo footprint, we now verify that the internal reddening derived from the extinction map discussed in §5.3 is very small and does not vary much spatially, using the the Draine et al. (2014) map which covers the full range of our UKIRT data in major axis, but not the full range in minor axis. However, the minor axis (northwest and southeast sections of M31’s halo) are expected to contain even less dust. The average extinction values derived from the corrected map discussed in §5.2.1 outside of a deprojected distance of 20 kpc are mag and mag, which are exactly equal to the foreground Milky Way reddening corrections also discussed in §5.2.1. For simplicity, we apply these small extinction corrections directly to our photometry before modeling the CMD, and then do not solve for extinction as part of the CMD modeling.
Figure 10 shows the NIR CMD of M31’s inner halo we used to measure its SFH ( kpc). The grey box indicates the selection region for AGB stars. Select PARSEC-COLIBRI isochrones (0.1, 1, 8 Gyr) are included for reference. There are AGB stars in this region. The average 50% completeness limit across the inner stellar halo is mag.
region 1 contains the Eastern Shelf, region 5 contains the Western Shelf, and region 4 contains NGC 205. Green polygons highlight the approximate positions of the the stellar structures which are a part of the GSS, and the pink ellipse corresponds to the dwarf galaxy NGC 205 in this map. The positions of the AGB stars and their region labels are available as Data behind the Figure.
@@ 5.3.2 Spatial Selection
The large number of AGB stars in M31’s inner halo allow us to measure a spatially resolved SFH. We divide our photometry into 6 spatial regions according to the following criteria. First, as illustrated in Figure 9, we prioritize placing three major substructures from McConnachie et al. (2018) into separate regions. Regions 1 and 5 contain the Eastern and Western Shelves, respectively, which have been identified as being radial “shell” features created from the Giant Stellar Stream (GSS) progenitor’s last pericentric passage (Ferguson et al., 2002, 2005; Fardal et al., 2006, 2008; Richardson et al., 2008; Fardal et al., 2012; Bernard et al., 2015b). Region 4 contains NGC 205 (M110), a satellite galaxy of M31. We note M32, another satellite galaxy of M31, overlaps on the sky with the inner disk of M31 where crowding was very high; the 50% completeness level in the J band across the area covered by M32 was J = 18.5 mag. Therefore, we were unable to probe the SFH of M32 using this photometry.
The remaining regions were chosen to have at least AGB stars, which is near the lower limit for the number of AGB stars needed for a robust SFH recovery (see Appendix D). The regions were delineated via a combination of radial and azimuthal lines, resulting in the irregular region shapes seen in Figure 9. The goal for these remaining sections is simply to provide as much resolved spatially information as possible, particularly in the radial direction, in light of other spatial and number of star constraints. The number of AGB stars ranged from AGB stars in region 1 to AGB stars in region 5.
@@ 5.3.3 Star Formation History
We run MATCH on each of the 6 spatial sub-regions. As discussed in Section 4, we measured the average stellar density of these 6 halo regions and used representative ASTs in three representative stellar density regions, as described in Table 2. We compute random uncertainties for each region following the HMC approach described in §3.
@@
Figure 11 and Figure 12 show the differential and cumulative SFHs for each region. The SFR density was calculated by normalizing the SFR by the projected area of each region. We note that there is fairly sizable variation in timing, duration, and amplitude of the SFHs as a function of position. Because the inner stellar halo of M31 is dominated by sub-structure, and not a clear smooth component, it is perhaps not surprising to see such variations in the SFHs and stellar mass growth as the inner halo was assembled from many individual pieces, each with their own formation histories.
In general, we find that most regions have prominent star formation Gyr ago and lower-level star formation Gyr ago. Many also exhibit some level of star formation at ages of Gyr. Expanding the NIR imaging and/or applying this method to other wide-field NIR imaging of M31 is an obvious path forward to exploring the SFH of its stellar halo at larger radii and to providing broader context with which to interpret our SFHs.
@@
In Figure 13, we plot the spatially-resolved SFH of M31’s inner stellar halo. Each of the 6 regions is color-coded by the best fit SFR at select lookback times. We note several interesting features. First, the clearest enhancements in SFR occurred Gyr ago. Second, Region 4, which includes NGC 205, shows prominent star formation and Gyr ago. The SFH of NGC 205 measured from deep HST optical imaging shows a similar enhancement in star formation activity Gyr ago (Savino et al., 2025). Older studies of NGC 205’s stellar populations qualitatively support the presence of intermediate age and even young stellar populations in NGC 205’s central region (e.g., Richer et al., 1984; Lee, 1996; Davidge, 2003, 2005; Rose et al., 2005; Sharina et al., 2006; Monaco et al., 2009). The interaction history of NGC 205 and M31 is not as well-studied as others in the M31 system, such as M32. There is a putative tidal ‘arc’ of NGC 205 that extends to the north ( McConnachie et al. 2004), though its association with NGC 205 is not definitive (e.g., Howley et al., 2008; Liu et al., 2024). An interesting future angle would be to combine wide-area kinematics, such as measured from recent work with DESI (Dey et al., 2023), and potentially Subaru/PFS in the future, with the AGB-star based SFHs to see if there are clear relationships between ages and kinematics, analogs to the disk heating study of Dorman et al. (2015). Finally, NGC 205 currently has no published proper motions. An orbital history, in tandem with our SFHs, would be useful in helping to decipher the interaction history between NGC 205 and M31.
Figure 14 shows the global SFH of M31’s inner halo (i.e., all 6 regions summed together), along with the corresponding 68% uncertainties. The overall shape of the SFH is similar to many of the individual regions, with a prominent burst of star formation Gyr ago, and steady and variable star formation from Gyr ago. Our global SFH shows that over the past Gyr, the inner stellar halo of M31 formed , % of which appears to have occurred Gyr ago.
@@
Figure 15 shows the ‘quenching’ time in each region. As a proxy for ‘quenching’ we use , the lookback time at which 90% of the stellar mass formed in each region’s cumulative SFH. The bulk of SF ceased earlier in regions with larger values of compared with regions with smaller values of .
@@ 6 Caveats & Limitations
First, in this paper, we only examined AGB star SFHs back to Gyr ago. This oldest age restriction was motivated by the conservative choice to empirically vet our data against another SFH and because of the limited photometry depth of this dataset. However, in principle, as discussed in §2, it is possible to measure an AGB star SFH back to the oldest epoch of star formation (i.e., Gyr ago). While these ancient stars are unlikely to be found in the TP-AGB phase, they will appear as E-AGB stars on a NIR CMD. Indeed, ancient MW globular clusters (GC) (e.g., NGC 6752; Norris et al. 1981) do have confirmed AGB stars, providing affirmation of this possibility. The limiting factor for testing this possibility is the lack of a suitable dataset (i.e., it must include the ancient MSTO and AGB stars). We plan to continue this investigation in future papers.
Second, as illustrated in Appendix D, a robust SFH requires several hundred AGB stars. We adopt a conservative recommendation of AGB stars to recover the most accurate and precise SFH.
@@
This minimum number is comfortably met by systems into the dwarf galaxy regime. For example, the SMC () and NGC 6822 (; McConnachie 2012) each contain AGB stars (e.g., Cioni & Habing, 2005; Boyer et al., 2011), while other fainter systems have several hundred AGB stars as well (e.g., Dalcanton et al., 2012a; Boyer et al., 2015). While galaxy luminosity provides a good rule of thumb for how many AGB stars to expect, the SFH is clearly an important component. Systems that are purely ancient will have fewer AGB than those with more recent star formation (i.e., within the last several Gyr). Even fairly luminous systems with ‘pauses’ in the SFHs at intermediate ages (e.g., WLM; Albers et al., 2019; McQuinn et al., 2024a) may have fewer AGB stars than their luminosities would suggest.
@@
Third, we did not calculate systematic uncertainties for our AGB star SFHs, due to the lack of multiple available AGB stellar models. Dolphin (2012) presents a means to estimate uncertainties on SFHs measured from resolved star CMDs that are designed to capture variations in the SFHs due to choice in stellar models. The method for computing these systematics relies on being able to measure SFHs on test datasets using different stellar models. Because COLIBRI uniquely provide suitable AGB star models, it is not possible to compute analogous systematic uncertainties for our AGB star SFHs. Even within the COLIBRI models, specific choices are made that can affect the age characteristics of the AGB stars. For example, our basis models for this analysis were generated with a Reimers mass loss value of for RGB stars. Large changes to this value can affect the evolution of AGB stars. For example, a much higher mass-loss rate on the RGB does not produce AGB stars at the oldest ages. It is not clear that such large mass-loss rates on the RGB are supported by data, but we nevertheless note this possibility. This type of choice is not unique to AGB star models. For example, the modeling of metal-diffusion in lower-mass stars can affect the lifetimes of MSTO stars, which can lead to systematically younger or older ages (e.g., Hidalgo et al., 2018).
@@
Finally, we note that only NIR data are suitable for AGB star SFHs. In the optical, there is too much overlap between the RGB and AGB to provide robust SFHs (e.g., see Figure 5). We have not conducted an exhaustive exploration of all possible filter combinations, but our findings thus far strongly suggest this approach is optimal with two NIR filters. Age sensitivity is improved when both sides of the typical AGB SED peak are comfortably sampled (e.g., one band at m and another at m). We also note that unfortunately the HST WFC3/IR F110WF160W color cannot be used for determining AGB star SFHs; SSPs simulated from the COLIBRI stellar models by Dalcanton et al. (2012a) indicated that separating carbon-rich and oxygen-rich AGB stars with F110WF160W colors is difficult because they occupy the same CMD space in that filter combination.
@@ 7 Future Prospects
The type of JWST dataset needed to measure AGB-star SFHs is well-matched to the JAGB (e.g., Lee et al. 2024; Li et al. 2024), and to some extent, TRGB distance programs (e.g., Anand et al. 2024a, b; Hoyt et al. 2024; Newman et al. 2024). The JAGB method requires at least a few hundred AGB stars in the J region of the NIR CMD and is often measured in filter combinations that sample either side for the SED peak of AGB stars. The high luminosity of AGB stars coupled with the characteristics of JWST suggest that AGB-based SFHs and JAGB distances already work out to 20 Mpc (e.g., Anand et al., 2024a, b; Hoyt et al., 2024; Lee et al., 2024) using hours of integration time. It is seems clear that JAGB distances and our SFH method can be extended to galaxies out to many tens of Mpc with suitably well-planned observations (e.g., to avoid too much crowding).
Programs that aim to measure TRGB distances with JWST may also be suitable for AGB-based SFHs. The filter combination provides for precise TRGBs, but is not particularly good for clearly separating all AGB and RGB stars in CMD space (see e.g., various CMDs in Anand et al. 2024a, b; Newman et al. 2024). However, the filter combination appears to enable precise (.5%) TRGB distances (Newman et al., 2024) and appears to be a good filter combination of AGB star SFHs as AGB stars are both brighter in these bands, have little overlap with the RGB, and sample both sides of the SEDs peak for most AGB stars. We illustrate this further in Appendix B.
Roman and Euclid offer similar potentials for AGB star SFHs, though their filter ranges are a little more limited than JWST. We have not explored the specific capabilities of these facilities, but based on the analysis in this paper, AGB star SFHs from their NIR CMDs appear promising as well.
@@
However, robust determinations of SFHs derived from AGB stars will require investigations into the full systematic uncertainties of the fits, similar to what has been done for SFHs based on optical CMDs (e.g., Weisz et al., 2011; Dolphin, 2012). As discussed in §6, we highly advocate for the development of an independent set of very detailed TP-AGB models. In addition to aiding AGB star SFH studies like this one and others described in §1, another set of AGB models would improve our knowledge of the contribution of AGB stars to the integrated SEDs of galaxies at all cosmic epochs (e.g., Melbourne et al., 2012) and help with the calibration of the AGB star extragalactic distance scale, which use the P-L relations of Mira variable stars (e.g., Huang et al. 2020, 2024) or the mean luminosities of carbon stars (JAGB method; e.g., Madore & Freedman 2020; Ripoche et al. 2020; Zgirski et al. 2021; Lee et al. 2024).